Hi, I am having trouble working out this question, help would be much appreciated.
An electrically charged particle generates an electric field around it. The strength, E, of the electric field (in NC^-1) is directly proportional to the size of the charge of the particle, Q (in Coulombs, C), and inversely proportional to the square of the distance from it, d in metres.
(i). Set up the proportionality relationship
This is what I think the formula is but not sure if it is correct?
E (NC^-1)= k * (Q/D^2)
And then theirs question (ii)
(ii). Find the constant of proportionality if the electric field strength 3.00 cm from a 3.5 uC charge is 3.5 x 10^7 NC^-1
This is what I've worked out so far not sure if it is correct though?
3.5 * 10^7 = K * Q/(0.03)^2
3.5 * 10^7 = K * Q/(0.09)

Answer by rothauserc(4718)

uC is 1 * 10^-6 Coulombs
:
3.5 uC = 3.5 * 10^-6 Coulombs = Q
:
3.5 * 10^7 = K * Q/(0.09)
:
(9 * 10^-2)(3.5 * 10^7) = K * Q
:
31.5 * 10^5 = K * (3.5 * 10^-6)
:
K = (31.5 * 10^5) / (3.5 * 10^-6)
:
K = 9 * 10^11